OpenStudy (anonymous):

if the function f is defined by the equation f(x,y)=x^2 y^3 and f(a,b)=10, what is the value of f(2a,2b)?

OpenStudy (raden):

f(a,b) = 10 ---> a^2b^3 = 10 f(2a,2b) = (2a)^2(2b)^3 = 4a^2*8b^3 = 32a^2b^3 now, u just subtitute the value of a^2b^3 = 10

OpenStudy (anonymous):

i didnt get u

OpenStudy (sirm3d):

the first substitution yields \[a^2b^3=10 \]

OpenStudy (sirm3d):

the 2nd evalution yields \[f(2a,2b)=(2a)^2(2b)^3\]

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

options 50,100,160,320

OpenStudy (sirm3d):

simplify (2a)^2(2b)^3 first

OpenStudy (anonymous):

4a^2 8a^3

OpenStudy (sirm3d):

you mean 4a^28b^3

OpenStudy (anonymous):

yup

OpenStudy (sirm3d):

or 32a^2b^3

OpenStudy (anonymous):

how 32?

OpenStudy (sirm3d):

product of the numerical coefficients

OpenStudy (anonymous):

show steps plz

OpenStudy (sirm3d):

\[(4a^2)(8b^3)=32a^2b^3\]

OpenStudy (anonymous):

now i got it :)

OpenStudy (sirm3d):

so the answer to the problem is?

OpenStudy (anonymous):

32 X 10

OpenStudy (anonymous):

320?

OpenStudy (sirm3d):

bingo

OpenStudy (anonymous):

what will we substitude in place of a n b?

OpenStudy (anonymous):

okaay got it

OpenStudy (anonymous):

thanks